OpenStudy (monkey*):

Will give medal! 2 Questions! x-2/x+4 = x-6/x-4 A) X=8 B) X=3 C)X=9 D)X=5

OpenStudy (monkey*):

I think the answer is a.

OpenStudy (lil_drew59):

b is correct let check just to make sure

OpenStudy (lil_drew59):

yea its x=3

OpenStudy (monkey*):

How did you get that?

OpenStudy (lil_drew59):

give me a sec n ill show you

OpenStudy (lil_drew59):

aye is that 8 -8

OpenStudy (lil_drew59):

Simplifying x + -2x + 4 = x + -6x + -4 Reorder the terms: 4 + x + -2x = x + -6x + -4 Combine like terms: x + -2x = -1x 4 + -1x = x + -6x + -4 Reorder the terms: 4 + -1x = -4 + x + -6x Combine like terms: x + -6x = -5x 4 + -1x = -4 + -5x Solving 4 + -1x = -4 + -5x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '5x' to each side of the equation. 4 + -1x + 5x = -4 + -5x + 5x Combine like terms: -1x + 5x = 4x 4 + 4x = -4 + -5x + 5x Combine like terms: -5x + 5x = 0 4 + 4x = -4 + 0 4 + 4x = -4 Add '-4' to each side of the equation. 4 + -4 + 4x = -4 + -4 Combine like terms: 4 + -4 = 0 0 + 4x = -4 + -4 4x = -4 + -4 Combine like terms: -4 + -4 = -8 4x = -8 Divide each side by '4'. x = -2 Simplifying x = -2

OpenStudy (lil_drew59):

i rounded tha 2 upward cause u didnt have that as a answer choice